# Single-token vs Two-token Blockchain Tokenomics

### v1

@@ Abstract
We consider long-term equilibria that arise in the tokenomics design of proof-of-stake (PoS) blockchain systems that comprise of users and validators, both striving to maximize their own utilities. Validators are system maintainers who get rewarded with tokens for performing the work necessary for the system to function properly, while users compete and pay with such tokens for getting a desired system service level. We study how the system service provision and suitable rewards schemes together can lead to equilibria with desirable characteristics (1) viability: the system keeps parties engaged, (2) decentralization: multiple validators are participating, (3) stability: the price path of the underlying token used to transact with the system does not change widely over time, and (4) feasibility: the mechanism is easy to implement as a smart contract, i.e., it does not require fiat reserves on-chain for buy back of tokens or to perform bookkeeping of exponentially growing token holdings. Furthermore, we consider both the common single-token PoS model and a less widely used two-token approach (that roughly, utilizes one token for the users to pay the transaction fees and a different token for the validators to participate in the PoS protocol and get rewarded). Our approach demonstrates, for the first time to our knowledge, concrete advantages of the two-token approach in terms of the ability of the mechanism to reach equilibrium.
@@ 1 Introduction
    Viability. The system that keeps all involved parties actively engaged: validators to guarantee the protocol being alive and securely running, and the users to guarantee enough fees are collected at all time.
@@
    Decentralization. A set of validators each equally engaged in the protocol, receiving adequate rewards, while also having significant stake in the system operation (e.g. in the case of proof-of-stake (PoS) systems, validators “staking” high enough amount of tokens).
@@
    Stability. Reasonably stable token prices, meaning that the price of the token required to issue transactions over time does not change or does not become inflationary (decreasing prices).
@@ 1.1 Main contributions
We give both boundary conditions and actual policies (mechanisms) that implement equilibria with all desired features: viability, decentralization, stability, and feasibility. Specifically, the resulting mechanism only needs to adjust the total rewards allocated to the validators in the next round, though it does so in a somewhat counterintuitive way:
If the demand for tokens drops, instead of buying back tokens, the mechanism increases the rewards allocated to validators.
This idea steams from our analysis of equilibria that maintain a given price path (e.g., stable prices):
@@
The main catch here is that these two quantities are not necessarily related to each other, and the system can actually adjust the rewards in order to match changes in the value of the service, and keep the desired prices without performing token buy backs. If, at any point in time, the value of the service level decreases, users will buy fewer tokens, thereby reducing demand and potentially necessitating token buybacks. Rather than engaging in token buybacks, the system can opt to increase the rewards allocated to validators. As validators compete for the rewards, they react by staking more tokens, which they must acquire from the market. This action effectively restores the demand for tokens to align with the supply and avoids buybacks.
  • Single Token Equilibria. We provide a mechanism which adjusts the rewards at every round based on the current rewards and the next round service value using a single token for accounting. In order for this mechanism to work, and keep the token rewards and staked holdings under control for feasibility, the mechanism needs to have some global information about the service value time series (e.g., a lower bound, cf. Theorem 2 and Corollary 2).
@@
Intuitively, the mechanism strives to use the largest possible rewards (so to guarantee decentralization and security – a good margin for the validators profit), while not exceeding at any point in time the value which triggers an uncontrolled growth in the rewards and in the staked amounts of tokens (Section 3.3.1). In other words, the mechanism provides the highest possible rewards (i.e., decentralization and security guarantee) while satisfying the other three desiderata (viable system, stable prices, no buy backs). In Section 3.4 we describe in detail the implementation of our mechanism for the case of stable prices. This implementation includes the strategies for users and validators, showing that all information needed is the next round service level offered by the system (despite their simplicity, these strategies still constitute an equilibrium in the underlining repeated game, i.e., when more complex time-dependent strategies are in principle possible).
@@
The above results apply to a single token setting which is the most common in PoS systems (e.g., Ethereum and Cardano operate in this way among the top cryptocurrencies). We next explore the setting of two token PoS in which one token is used to get the service and another token is used for staking (and governance). While less popular, this type of setup has been adopted in the Neo cryptocurrency [^1] See https://docs.neo.org/docs/en-us/index.html. and can be implemented in other systems as well (e.g., in the Cosmos SDK [^2] See https://docs.cosmos.network/main/learn/beginner/gas-fees.). To investigate the potential of these mechanisms, we introduce a natural variant of our model with two tokens. In order to maintain feasibility and avoid buy backs, while maintaining desirable prices, we identify the necessary conditions that equilibria (mechanism) must guarantee in terms of rewards:
@@
  • Double Token Equilibira. The two token mechanism is presented in Section 4.3. The main advantage of the resulting rewards formula is that it is simpler than in the single token setting and the mechanism does not need to know “global information” about the service value time series — merely a one step lookahead suffices (cf. Theorem 4). As a result, the mechanism achieves our objectives with less information about the service value time series.
@@
Intuitively, the advantage of “decoupling” the users from the validators, using different tokens, allows to better react and absorb shocks and fluctuations in the value of the service while maintaining the price of the users’ token stable and keeping all our feasibility desiderata. The price of the token used for staking increase “gradually”, while the amount of staked tokens remains constant, without scarifying incentive compatibility. Hence, the security of the protocol is preserved (the value of the staked tokens increases). This should be compared and contrasted with the single-token mechanism, which either (i) needs to know a global property of the service level time series, or in any case, be informed well in advance about a sudden drop in service value in order to be able to gradually decrease the rewards (as it can only adjust the rewards according to a certain “decreasing curve”), or (ii) it should proactively keep rewards well below the possible value (thus achieving a suboptimal security level for the protocol).
@@
Related work. In the single-token setting, our results build on the model of Häfner (2023). Compared to that prior work, our results accommodate a more general class of equilibria with respect to their features, that facilitate stable prices without the requirement to employ buy backs. To the best of our knowledge, ours is the first analytical model for such systems to study their equilibria and related token price paths achieving all four desiderata. Prior work includes the effects of tokens regarding user adoption and equilibrium section Bakos and Halaburda (2019, 2022); Sockin and Xiong (2023); Li and Mann (2018), or using token supply and buyback to promote optimal growth and service provision Cong et al. (2021, 2022). Different aspects of the monetary characteristics of tokens have been studied in Schilling and Uhlig (2019); Pagnotta (2022); Prat et al. (2021), while Kiayias et al. (2023) considered token burning. Typically the users receive the most modelling attention, but Chitra (2021) expand on the strategy space of validators and their alternative uses for tokens. While all previous work has focused on the single token design, Dimitri (2023) has considered a two-token model.
@@ 2 Model description and notation (single token)
The condition that tokens holding must be strictly positive at all time steps is identical to the definition of equilibrium in Häfner (2023). In particular, it implies that the selling/buying strategies at such an equilibrium satisfy the selling constraints (6) and (5) with a strict inequality.
@@
Lemma 1.
(Häfner, 2023, Lemma 1) For the following special case of reward sharing and service schemes,
@@ 3.2 Generic symmetric equilibria
In the following we focus on the interesting case of prices having a constant multiplicative growth, that is, for all $t\geq 1$
@@
thus implying that it must hold $\kappa_{R}>0$ and $\kappa_{S}>0$.
@@ 3.3 No buy back condition and its implications
Like in the original model by Häfner (2023), in the model considered so far the system is allowed to buy or to sell the additional tokens required during each subround II to match demand with supply. In this section, we refine the concept of equilibrium, by requiring that the system does not have to buy back tokens (and perhaps does not sell tokens either if demand and suppy match at equilibrium).
@@
We say that a symmetric equilibrium satisfies the no buy back condition if no additional tokens are bought at any round by the system, that is, $mb_{U}^{(t)}\geq ns_{V}^{(t)}$ for all $t$.
Intuitively speaking, the following theorem says that increasing the rewards at the next round does help for the no buy back condition. The intuitive, though perhaps surprising, reason is that this will make rewards more attractive for the validators who then stake more tokens (and thus sell less on the market).
@@
$mb_{U}^{(t)}=m\mathbb{TK}_{U}^{(t+1)}=\frac{S^{(t+1)}}{\mathbb{PRICE}^{(t+1)}\cdot Ser2Fees}=\frac{S^{(t+1)}\cdot(\delta\mathcal{R})^{t+1}}{\mathbb{PRICE}^{(0)}\cdot Ser2Fees}$
@@
The above theorem provides a recursive (algorithmic) formula for the rewards, which leads to our mechanism described in Section 3.4 below. The same theorem also implies the following bound on the growth of the rewards necessary to guarantee the no buy back.
@@
The minimal rewards satisfying the no buy back condition tightly are equal to
@@ 3.4 Policy that accommodates and implements the stable price equilibrium
In this section, we consider the fundamental question of how one can compute and implement an equilibrium with stable prices satisfying the no buy back condition. Stable prices require to keep a precise amount of tokens of users (on one side) and of tokens of validators (on the other side) given by Corollary 1. The no buy back condition requires the system to set the rewards so to satisfy the condition in Theorem 2. For the sake of exposition, we next describe an implementation for the reward and service fee schemes in (16).
@@
      Set the next service level $S^{(t+1)}$ and the next total rewards $R^{(t+1)}$ so that
@@
      Announce the next service level $S^{(t+1)}$ and the next rewards $R^{(t+1)}$ before the token buy or sell (Subround II) begins.
@@ 4 A two-token model
System token $\mathbf{A}$ purpose token $\mathbf{B}$ purpose reference
DFINITY ICP staking & rewards Cycle computation (service) (Team, 2022)
NEO NEO staking GAS pay transactions (neo.org, )
Axie Infinity AXS governance and staking SLP play (breeding Axie) (Axi, 2021)
@@ 4.3 Stable price mechanism in two-token model
We next‘ describe a simple mechanism which implements stable prices for token $\mathbf{B}$, while the price of $\mathbf{A}$ is strictly decreasing:

### v2

@@ Abstract
We consider long-term equilibria that arise in the tokenomics design of proof-of-stake (PoS) blockchain systems that comprise of users and validators, both striving to maximize their own utilities. Validators are system maintainers who get rewarded with tokens for performing the work necessary for the system to function properly, while users compete and pay with such tokens for getting a desired portion of the system service.
We study how the system service provision and suitable rewards schemes together can lead to equilibria with desirable characteristics (1) viability: the system keeps parties engaged, (2) decentralization: multiple validators are participating, (3) stability: the price path of the underlying token used to transact with the system does not change widely over time, and (4) feasibility: the mechanism is easy to implement as a smart contract, i.e., it does not require fiat reserves on-chain for buy back of tokens or to perform bookkeeping of exponentially growing token holdings. Our analysis enables to put forward a novel generic mechanism for blockchain “monetary policy” that we call quantitative rewarding (QR). We investigate how to implement QR in single-token and two-token proof of stake (PoS) blockchain systems. The latter are systems that utilize one token for the users to pay the transaction fees and a different token for the validators to participate in the PoS protocol and get rewarded. Our approach demonstrates a concrete advantage of the two-token setting in terms of the ability of the QR mechanism to be realized effectively and provide good equilibria. Our analysis also reveals an inherent limitation of the single token setting in terms of implementing an effective blockchain monetary policy — a distinction that is, to the best of our knowledge, highlighted for the first time.
@@ 1 Introduction
    Viability. The system keeps all involved parties actively engaged: validators to guarantee the protocol being alive and securely running, and the users to guarantee enough fees are collected at all time.
@@
    Decentralization and “Skin-in-the-game.” A set of validators each equally engaged in the protocol, receiving adequate rewards, while also having significant stake in the system operation (e.g. in the case of proof-of-stake (PoS) systems, validators “staking” high enough amount of tokens).
@@
    Stability. Reasonably stable token prices, meaning that the price of the token required to issue transactions over time does not change or becomes inflationary (decreasing prices).
@@ 1.1 Main contributions
We give both boundary conditions and actual policies (mechanisms) that implement equilibria with all desired features: viability, decentralization, stability, and feasibility. Specifically, our analysis leads us to put forward a novel mechanism for blockchain “monetary policy”:
Quantitative Rewarding (QR). The amount of rewards the system gives to the validators is suitably adjusted at each step to anticipate changes in the value of service.
This idea stems from our analysis of equilibria that maintain a given price path (e.g., stable prices):
@@
The main catch here is that these two quantities are not necessarily related to each other, and the system can actually adjust the rewards in order to match changes in the value of the service, and keep the desired prices without directly participating in the token market (e.g., as it would be the case via buy-backs). Observe that if, at any point in time, the value of the service level decreases, users will buy fewer tokens, thereby reducing demand and potentially necessitating token buybacks if one is to keep prices stable. Rather than engaging in token buybacks, QR increases the rewards allocated to validators. As validators compete for the rewards, they react by staking more tokens, which they must be acquired from the market. This action effectively restores the demand for tokens to align with the supply. The reverse effect takes place when the service level increases.
  • Single Token Equilibria. We investigate how to implement QR via a mechanism which adjusts the rewards at every round based on the current rewards and the next round service value using a single token for accounting. An important challenge that arises in this setting is that the rewards formula is subject to a recursive condition (see the lower bound in Theorem 2 and Corollary 2) which makes QR subject to a “feedback loop” effect that can lead to rewards explosion.
@@
As a result, while maintaining certain (minimal) rewards to ensure decentralization and security, it is possible that the system can trigger an uncontrolled growth in the amount of rewards and in the staked amounts of tokens (Section 3.3.1). In Section 3.4 we describe in detail the implementation of our mechanism for the case of stable prices. This implementation includes the strategies for users and validators, showing that all information needed is the next round service level offered by the system (despite their simplicity, these strategies still constitute an equilibrium in the underlining repeated game, i.e., when more complex time-dependent strategies are in principle possible).
@@
The above results apply to a single token setting which is the most common in PoS systems (e.g., Ethereum and Cardano operate in this way, for example, as PoS cryptocurrencies). We next explore the implementation of QR in two token PoS systems in which one token is used to get the service and another token is used for staking (and governance). While less common, this type of setup has been adopted in a few occasions: e.g., in the Neo cryptocurrency [^1] See https://docs.neo.org/docs/en-us/index.html. and can be implemented in other systems as well (e.g., in the Cosmos SDK [^2] See https://docs.cosmos.network/main/learn/beginner/gas-fees.). To investigate the potential of QR in this setting, we introduce a natural variant of our model with two tokens. In order to maintain feasibility and avoid buy backs, while maintaining desirable prices, we identify the necessary conditions that equilibria (mechanism) must guarantee in terms of rewards:
@@
  • Double Token Equilibria. The two token mechanism for QR is presented in Section 4.3. The main advantage of the resulting rewards formula is that it is simpler than in the single token setting and the mechanism does not need to know “global information” about the service value time series — merely a one step lookahead suffices (cf. Theorem 4). As an important corollary, the mechanism achieves our objectives and is capable of avoiding the feedback loop effect that takes place in the single token case that leads to an exponential explosion for rewards.
@@
Intuitively, the advantage of “decoupling” the users from the validators, using different tokens, allows to better react and absorb shocks and fluctuations in the value of the service while maintaining the price of the users’ token stable and keeping all our feasibility desiderata. The price of the token used for staking increases “gradually”, while the amount of staked tokens remains constant, without scarifying incentive compatibility. This should be compared and contrasted with the single-token setting, which may enter a feedback loop and require an increase of the validator rewards indefinitely along an exponential curve (even e.g., after the service time series stabilizes) — something that appears to be an inherent distinction between the single token and the two token setting in our model.
@@
Related work. In the single-token setting, our results build on the model of Häfner (2023). Compared to that prior work, our results accommodate a more general class of equilibria with respect to their features, that facilitate stable prices without the requirement to employ buy backs. To the best of our knowledge, ours is the first analytical model for such systems to study their equilibria and related token price paths achieving all four desiderata. Prior work includes the effects of tokens regarding user adoption and equilibrium section Bakos and Halaburda (2019, 2022); Li and Mann (2018); Sockin and Xiong (2023), or using token supply and buyback to promote optimal growth and service provision Cong et al. (2021, 2022b). Different aspects of the monetary characteristics of tokens have been studied in Schilling and Uhlig (2019); Pagnotta (2022); Prat et al. (2021), while Kiayias et al. (2023) considered token burning. Typically the users receive the most modelling attention, but Chitra (2021) expand on the strategy space of validators and their alternative uses for tokens. Further works modelling validators staking strategies are Cong et al. (2022a); Jermann (2023); Saleh (2021); John et al. (2021). While all previous work has focused on the single token design, Dimitri (2023) has considered a two-token model.
@@ 2 Model description and notation (single token)
The condition that tokens holding must be strictly positive at all time steps is identical to the definition of equilibrium in Häfner (2023). In particular, it implies that the selling/buying strategies at such an equilibrium satisfy the selling constraints (6) and (5) with a strict inequality. This condition captures our viability requirement of keeping all parties engaged.
@@
Lemma 1 (Lemma 1 in Häfner (2023)).
For the following special case of reward sharing and service schemes,
@@ 3.2 Generic symmetric equilibria
In the following we focus on the interesting case of prices having a constant multiplicative growth, which includes stable prices as a special case. That is, for all $t\geq 1$
@@
thus implying that it must hold $\kappa_{R}>0$ and $\kappa_{S}>0$. Theorem 1 then implies the following.
@@ 3.3 No buy back condition and its implications
In the model considered so far, the system is allowed to buy or to sell the additional tokens required during each subround II to match demand with supply. In this section, we refine a concept of equilibrium, by requiring that the system does not have to buy back tokens (and perhaps does not sell tokens either if demand and supply match at equilibrium). We stress that this refers only to the “spot market” subround ĨI where tokens are exchanged for money. Indeed, these “Lagos-Wright” two-stage models have a simple implementation (subround I): The system either burns some fees (when rewards are less than the paid fees) or mints additional new tokens for rewards (when rewards are more than paid fees); observe that no monetary reserve is needed for these operations.
@@
We say that a symmetric equilibrium satisfies the no buy back condition if no additional tokens are bought at any round by the system, that is, $mb_{U}^{(t)}\geq ns_{V}^{(t)}$ for all $t$. Additionally, we say that the no buy back condition holds tightly if demand matches supply, that is, $mb_{U}^{(t)}=ns_{V}^{(t)}$.
Note that the quantities $m\cdot b_{U}$ and $n\cdot s_{V}$ correspond to the token demand and supply in the “market” subround II, respectively. As we formally prove below, the demand and supply are determined by the service level and rewards, respectively:
  • Demand ($mb_{U}^{(t)}$ from users) decreases if the next round’s service level decreases (Equation (30) below).
  • Supply ($ns_{V}^{(t)}$ from validators) decreases with the next round’s rewards (Equation (31) below).
This naturally suggests the following quantitative rewarding approach to avoid buybacks:
  1. 1.
    (demand decreases $\Rightarrow$ increase rewards.) Whenever the next round’s service level “decreases significantly,” in order to adjust supply to the decreased demand, we increase the next round’s rewards. This will make rewards more attractive for the validators, who then stake more tokens (and thus sell less on the market).
  2. 2.
    (demand increases $\Rightarrow$ decrease rewards.) Whenever the next round’s service level “increases significantly,” we can decrease the next round’s rewards. This will have the opposite effect on validators, who will sell more tokens on the market.
The first case is necessary to avoid buybacks (satisfy Definition 3). The second operation turns out to be useful for (i) restoring payments to a lower value when possible and (ii) satisfying the no buybacks condition tightly, meaning that demand never exceeds supply ($mb_{U}=ns_{V}$). This stronger condition has the advantage that the system does not even need to implement a minting mechanism for users to buy tokens (see Remark 4 below).
The following theorem establishes an equivalence between the no buybacks condition and the above quantitative rewarding (rewards increase). In particular, it quantifies the minimum rewards for which it is possible to avoid buybacks at a given round.
@@
$mb_{U}^{(t)}=m\mathbb{TK}_{U}^{(t+1)}=\frac{S^{(t+1)}}{\mathbb{PRICE}^{(t+1)}\cdot Ser2Fees}=\frac{S^{(t+1)}\cdot(\delta\mathcal{R})^{t+1}}{\mathbb{PRICE}^{(0)}\cdot Ser2Fees}$ (30)
@@
$ns_{V}^{(t)}$ $=n\left(\mathbb{TK}_{V}^{(t)}-\mathbb{TK}_{V}^{(t+1)}+R^{(t)}\cdot r(1/n)\right)=\frac{R^{(t)}}{Rew2Stake}-\frac{R^{(t+1)}}{Rew2Stake}+R^{(t)}\cdot n\cdot r(1/n)\ .$ (31)
@@
The above theorem provides a recursive (algorithmic) formula for the rewards, which leads to our mechanism described in Section 3.4 below. By unfolding the recursion (28) in same theorem, we get the following bound on the growth of the rewards necessary to guarantee the no buy back.
@@
The minimal rewards satisfying the no buy back condition (tightly) are equal to
@@
We stress that the minimal rewards satisfying the no buy back condition must satisfy the no buy back condition tightly, and thus must coincide with the rewards $R_{single}^{(t)}$ in Corollary 2.
Remark 4 (avoid users minting with tight no buy back).
Observe that the no buy back condition only requires supply to not exceed demand (Definition 3). Hence, it is possible that (users’) demand exceeds (validators’) supply ($mb_{U}^{(t)}-ns_{V}^{(t)}>0$). In this case, the system has to provide this difference of tokens to the market (users). On the one hand, it allows the system to act as a net seller and to make some profit (like in some cases in Häfner (2023)). On the other hand, there are scenarios in which this might be difficult or undesirable (typically, it requires the system to implement a “special” minting mechanism that allows users to “buy” tokens). In such cases, we require a tight version of the no buy back condition (Definition 3) in which demand matches supply. The analysis of this tight no buy back condition is given by Corollary 2, while the formula in Theorem 2 with equality gives the recursion for the rewards.
As we discuss below, implementing the quantitative rewarding can be difficult due to the possibility of an “uncontrolled” rewards growth. Indeed, Corollary 2 states that the initial rewards $R^{(0)}$ cannot exceed a certain value unless rewards grow exponentially. This value, critically, depends on future service levels. In Section 3.4, we describe an implementation under the assumption that the initial rewards $R^{(0)}$ can be set properly by the system (either having full knowledge about future service levels, a good estimate, or some partial control over them).
@@ 3.3.1 Main dilemma: Uncontrolled growth of rewards
The above uncontrolled growth represents a potential “death spiral” stemming from necessary equilibrium conditions. First, it leads to impractical implementation due to numerical explosion in ledger validators accounts (thus violating feasibility). Second, it may even render such equilibria nonexistent for some parameter combinations. Note that adapting the rewards requires knowledge about shocks far in the future. Indeed, Theorem 2 states that rewards cannot be reduced arbitrarily from one step to the next. Such powerful oracles are generally not available, and solutions based on them also violate feasibility.
@@ 3.4 Policy that accommodates and implements the stable price equilibrium
In this section, we consider the fundamental question of how one can compute and implement an equilibrium with stable prices satisfying the no buy back condition. Stable prices require to keep a precise amount of tokens of users (on one side) and of tokens of validators (on the other side) given by Corollary 1. The no buy back condition requires the system to set the rewards so to satisfy the condition in Theorem 2. For the sake of exposition, we next describe an implementation for the reward and service fee schemes in (16). We should point out that this implementation does not specify how to set the initial rewards $R^{(0)}$ which is the crucial point to avoid rewards explosion (Corollary 2).
@@
      Given the next service level $S^{(t+1)}$, set the next total rewards $R^{(t+1)}$ so that
@@
      Announce the next rewards $R^{(t+1)}$ before the token buy or sell (Subround II) begins.
@@ 4 A two-token model
System token $\mathbf{A}$ purpose token $\mathbf{B}$ purpose
DFINITY (Team, 2022) ICP staking & rewards Cycle computation (service)
NEO (neo.org, ) NEO staking GAS pay transactions
Axie Infinity (Axi, 2021) AXS governance and staking SLP play (breeding Axie)
@@ 4.2 No buy back in the two-token model
The following theorem provides a bound on the rewards that depends only on the next round’s service level. The main advantage of the two-token model is that rewards can be adjusted “immediately” to the service level fluctuations, and they need to be low only when the latter is low. The mechanism in Section 4.3 implements this idea.
@@ 4.3 Stable price mechanism in two-token model
We next describe a simple mechanism which implements stable prices for token $\mathbf{B}$, while the price of $\mathbf{A}$ is strictly decreasing: